(no subject)

John Shields (jshields@ksera)
Wed, 2 Jun 1999 15:31:52 -0500

Shift Summary for Day shift, 06/02/99

On Shift: John Shields
Pat Toale

Others Present: Jason 1&2, Rick K., Ed, Sydney, S. Laplace, Sasha, etc.

Hours of beam delivered at all intensities: 6.0
Hours of beam delivered at nominal intensity: 6.0
Hours of 99 data (special runs not included): 6.0
Hours of special runs (laser scans, muon runs...): 0.8 hrs

Accounting of lost time:

tevatron down, controlled access

========================================================================
Runs ended this shift:

Run 13290 - 832 ECAL
Run 13291 - 832 ECAL
Run 13292 - 832 ECAL
Run 13293 - 832 ECAL
Run 13294 - 832 ECAL
Run 13295 - 832 ECAL
Run 13296 - 832 Laser Scan

Tapes brought to FCC this shift: 0

======================================================================

Accesses during shift: 1

=====================================================================

Description of Shift:

7:15 Gateway dies after one spill. This hanging NOT correlated to
stopping the trigger

8:30 DC histograms look lousy. Probably since only one booster out
of 12 used --> instantanious intensity ~1E13!!

11:00 Margaret Votavez & Vivian look into Gateway problems.

11:30 Gateway problems apparently fixed after M & V increase number
of event buffers

13:05 Val notes that the tracking timing may be slightly off, resulting
in event hits that are apparently AFTER spill has terminated.

13:15 beam dies. bummer.

13:30 run control CONTINUES to hang up when STOP run-- it's been doing
this pretty much all day.

13:45 P. Shanahan requests long laser scan. Shift crew magnanimously
complies.

13:40 called PREP about pink monitors. *Naturally* PREP said they had
no replacements, so a service call was placed instead.

14:30 New E/P constants installed: CSQ6
validity range 13285-99999

14:50 SL does offline electron calib. run, using run #13285 data
It turns out the channels w. the least electrons have 10 electrons,
indicating there are enough e- to do a fair calibration.

15:20 Rick K, Ed B, Hogan, Jim G, Jason L, & Elizabeth go on controlled
access to remove regenerator, fix filter wheel on laser, etc.